The output voltage of an AC generator is given by Av = (150 V)sin(70nt). The generator is connected across a 25.0-12 resistorWill rate for answer with work!
Answer #1

a) Vo = 150 V

b)

w = 70 pi

2 pi f = 70 pi

f = 35 Hz

c)

Vr = 150 / sqrt(2) = 106 V

d)

ir = Vr / R = 106 / 25 = 4.2426 A

e)

io = 150 / 25 = 6 A

f)

P = Vr ir = 106* 4.2426 = 450 W

g)

i = io sin (70 pi t)

i = 6* sin ( 70 pi * 0.005)

i = 5.346 A

h)

option (b) is correct

=====

Comment before rate in case any doubt, will reply for sure.. goodluck